10. D&G Enterprises issues bonds with a $1,000 face value that make coupon payments of $40 every 3 months. What is the coupon rate?

4. You are supposed to receive $3,000 four years from now. At an interest rate of 7%, what is that $3,000 worth today?

A.

$2,288.69

B.

$2,931.01

C.

$2,205.09

D.

$2,537.08

E.

$2,699.01
Answer #1

Coupon rate = coupon rate = 16.00% (40*4)/1000]Worth today = pv = future value / (1+r)^t $2,288.69 3,000 = (1+0.07)-4)